OpenStudy (anonymous):

Find the general solution of x'=(2, 3, -1, -2)x+(e^t, t).

OpenStudy (loser66):

hey, girl, I really don't know how to solve it. @ybarrap

OpenStudy (ybarrap):

I don't really have a clue on this one, maybe it's the format. Looks like a vector, but then again not.

OpenStudy (anonymous):

it's 2x2 matrix, 2 and 3 on the left, -1 and -2 on the right. e^t on top, t on bottom. but I know that the general solution for 2x2 matrix, I got the answer but not the other part.

OpenStudy (anonymous):

\[x'=\begin{pmatrix}2&-1\\3&-2\end{pmatrix}x+\begin{pmatrix}e^t\\t\end{pmatrix}~~?\] I suggest undetermined coefficients, but only because I don't remember how to use the other methods... Do you need help with finding the eigenvalues, or the nonhomogeneous solution as well?

OpenStudy (anonymous):

Anyway, here's the setup for the eigenvalues: \[\begin{vmatrix}2-\lambda&-1\\3&-2-\lambda\end{vmatrix}=-(4-\lambda^2)+3=0~~\Rightarrow~~\lambda_1=1,\lambda_2=-1\] Let \(\vec{\eta}_1\) be the associated eigenvector for \(\lambda_1\): \[\begin{pmatrix}1&-1\\3&-3\end{pmatrix}\begin{pmatrix}\eta_1\\\eta_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\] So you have \(\eta_1=\eta_2\). Simplest value to pick would be 1, so \(\vec{\eta}_1=\begin{pmatrix}1\\1\end{pmatrix}\).

OpenStudy (anonymous):

Let \(\vec{\eta}_2\) be the eigenvector for \(\lambda_2\): \[\begin{pmatrix}3&-1\\3&-1\end{pmatrix}\begin{pmatrix}\eta_1\\\eta_2\end{pmatrix}=\begin{pmatrix}0\\0\end{pmatrix}\] So you have \(3\eta_1=\eta_2\). Letting \(\eta_1=1\) gives you \(\eta_2=3\), so your other eigenvector is \(\vec{\eta}_2=\begin{pmatrix}1\\3\end{pmatrix}\).

OpenStudy (anonymous):

That should take care of the homogeneous solution. It's been a while since I've learned this stuff, but from what I remember, you use a guess of the form \[\vec{x_p}=e^t\vec{A}+t\vec{B}+\vec{C}\] or perhaps \[\vec{x_p}=te^t\vec{A}+t\vec{B}+\vec{C}\] or maybe even \[\vec{x_p}=e^t\vec{A}+te^t\vec{B}+t\vec{C}+\vec{D}\] As to which one is correct, I'm not sure.

OpenStudy (loser66):

As @SithsAndGiggles stated, the homogeneous part is \[x_h = C_1 e^t\left(\begin{matrix}1\\1\end {matrix}\right) + C_2e^{-t}\left(\begin{matrix}1\\3\end{matrix}\right)\] your g(t) is \(g(t) = \left[\begin{matrix}e^t\\t\end{matrix}\right]\) for non homogeneous part, you have to construct \(\psi (t)\) from the homogenous part , just take out the C and put e^t , e^-t inside. \[\psi(t) = \left[\begin{matrix}e^t& e^{-t}\\e^t &3e^{-t}\end{matrix}\right]\] \[\psi (t)^{-1}= \left[\begin{matrix}3e^{-t}&-e^{-t}\\-e^t&e^t\end{matrix}\right]\], replace t by s, you have \(\psi(s)\) then apply formula to calculate the nonhomogeneous solution, say it is \(x_p\) \[x_p= \psi(t)\int_0^{t }\psi (s)^{-1}g(s)ds\]

OpenStudy (loser66):

I calculate the integrand first, then put it under the integral, then time it with \(\psi (t)\) integrand: \(\huge\psi (s)^{-1}*g(s) = \left[\begin{matrix}3e^{-s}&-e^{-s}\\-e^s&e^s\end{matrix}\right]\left[\begin{matrix}e^s\\s\end{matrix}\right]\) \[=\huge\left[\begin{matrix}3- s e^{-s} \\e^s(s-e^s)\end{matrix}\right]\]

OpenStudy (loser66):

then, take integral of it \[\huge \int_0^{t}\left[\begin{matrix}3-se^{-s}\\e^s(s-e^s)\end{matrix}\right]ds= \left[\begin{matrix}e^{-t}(t+1)\\te^t-\frac{e^{2t}}{2}-e^t+\frac{3}{2}\end{matrix}\right]\]

OpenStudy (loser66):

wow, i must love you so much to do a long thing like this , ha!!! to me, you are that special. ??? !!!

OpenStudy (loser66):

Now, time it with \(\psi (t)\) don't mess up the order \[ \left[\begin{matrix}e^t& e^{-t}\\e^t &3e^{-t}\end{matrix}\right]\left[\begin{matrix}e^{-t}(t+1)\\te^t-\frac{e^{2t}}{2}-e^t+\frac{3}{2}\end{matrix}\right]\] you have the \(x_p\) then you must add homogenous part and this part to get the final answer.

OpenStudy (loser66):

sorry girl, I have piano class at 12, must practice before class. cannot give you more. but I think you can finish it,right?

OpenStudy (ybarrap):

Nice job @SithsAndGiggles and @Loser66 . @Idealist, here is a nice explanation of the technique they used here: http://tutorial.math.lamar.edu/Classes/DE/NonhomogeneousSystems.aspx

OpenStudy (ybarrap):

Don't forget, once you have @SithsAndGiggles complementary solution \(x_h\) and @Loser66 particular solution \(x_p\), combine to get general solution \(x(t)=x_h+x_p\).

OpenStudy (loser66):

what is your question @Idealist

OpenStudy (anonymous):

How did you get v(t)^-1?

OpenStudy (loser66):

it is inverse of \(\psi (t\) it's 2x2 matrix, quite easy to find it inverse, right?

OpenStudy (anonymous):

What's the symbol v(t) called?

OpenStudy (loser66):

psi

OpenStudy (anonymous):

Bye.

OpenStudy (anonymous):

But I got a wrong answer for the inverse. How did you get yours?

OpenStudy (loser66):

oh yeah, you are right, I have to divide it by 1/2

OpenStudy (loser66):

I am sorry girl.

OpenStudy (anonymous):

It's okay.

OpenStudy (anonymous):

How did you multiply to g(s) and got the answer?

OpenStudy (loser66):

I don't get your question.

OpenStudy (anonymous):

Like how did you multiply v(t)^-1 and g(s)

OpenStudy (loser66):

how to multiply a matrix with a matrix, right?

OpenStudy (anonymous):

Yes.

OpenStudy (loser66):